# Two-stage Cytopathological Image Synthesis for Augmenting Cervical Abnormality Screening

### 22 Feb 2024 v1

@@ Abstract
Automatic thin-prep cytologic test (TCT) screening can assist pathologists in finding cervical abnormality towards accurate and efficient cervical cancer diagnosis. Current automatic TCT screening systems mostly involve abnormal cervical cell detection, which generally requires large-scale and diverse training data with high-quality annotations to achieve promising performance. Pathological image synthesis is naturally raised to minimize the efforts in data collection and annotation. However, it is challenging to generate realistic large-size cytopathological images while simultaneously synthesizing visually plausible appearances for small-size abnormal cervical cells. In this paper, we propose a two-stage image synthesis framework to create synthetic data for augmenting cervical abnormality screening. In the first Global Image Generation stage, a Normal Image Generator is designed to generate cytopathological images full of normal cervical cells. In the second Local Cell Editing stage, normal cells are randomly selected from the generated images and then are converted to different types of abnormal cells using the proposed Abnormal Cell Synthesizer. Both Normal Image Generator and Abnormal Cell Synthesizer are built upon the pre-trained Stable Diffusion via parameter-efficient fine-tuning methods for customizing cytopathological image contents and extending spatial layout controllability, respectively. Our experiments demonstrate the synthetic image quality, diversity, and controllability of the proposed synthesis framework, and validate its data augmentation effectiveness in enhancing the performance of abnormal cervical cell detection.
@@ I Introduction
Recently, diffusion models have gained enormous popularity due to their remarkable capabilities to generate high-quality images [13, 14]. Learning from large-scale data, text-to-image (T2I) diffusion models, especially Stable Diffusion [15], have demonstrated unprecedented ability in creating a wide variety of image contents that closely resemble the input text descriptions. To unleash the potential of T2I diffusion models, many studies have explored parameter-efficient fine-tuning (PEFT) methods in pursuit of model customization [16, 17, 18, 19] or additional controllability [20, 21, 22]. The impressive progress has inspired us to employ PEFT techniques to tailor T2I diffusion models for both few-shot synthesis and fine-grained conditioning in cervical cytopathological image synthesis.
In this paper, we propose a novel cytopathological image synthesis framework in support of augmenting cervical abnormality screening. The proposed framework can synthesize diverse cytopathological images that incorporate cervical cells of various abnormality degrees as shown in Fig. 1. The synthesis process naturally brings in a bounding box when placing an abnormal cell in the generated image. The synthetic data can thus be used to train a detector for suspicious abnormal cervical cells, which is essential to clinical usage. We empirically set the synthetic image size to be $1024\times 1024$ with a microscope magnification of ×$\times 20$, which strikes a balance between the number of cells in each image and the content of the image that a detector can effectively handle. As illustrated in Fig. 2, the overall framework is decomposed into two stages below:
@@
  • Global Image Generation. Based on a hypothesis that the neighboring effects between cervical cells are eliminated after oscillation and centrifugation in WSI preparation, we propose to synthesize abnormal cells by individually editing the cell types. Specifically, we employ a simple in-house cell detector to localize the synthetic NILM cells. After randomly selecting a certain number of them, we apply the Abnormal Cell Synthesizer. It can effectively extend the conditioning functionality of the pre-trained Stable Diffusion and translate the selected NILM cells into abnormal cells of user-defined types.
@@
  • We incorporate PEFT techniques in building both Normal Image Generator and Abnormal Cell Synthesizer upon Stable Diffusion, the state-of-the-art pre-trained T2I model, thus significantly alleviating the need for large-scale data in training reliable generative models.
@@ III Method
In this paper, we propose to disentangle the overall cervical cytopathological image synthesis into two stages, namely Global Image Generation and Local Cell Editing. As demonstrated in Fig. 2, the first stage produces high-resolution normal cytopathological images, and the second stage synthesizes high-fidelity abnormal cell patches. In the next, Section III-A reviews preliminary information about Stable Diffusion, which serves the pre-trained foundation in our work. Section III-B and Section III-C present Global Image Generation and Local Cell Editing in detail, respectively.
@@ V Conclusion and Discussion
Cytopathological image synthesis can minimize the efforts in data collection and annotation for cervical abnormality screening. However, it is difficult to generate a large-size cytopathological image while synthesizing visually realistic cellular features for small-size abnormal cells. To overcome this challenge, we propose a novel two-stage cervical cytopathological image synthesis framework based on diffusion models for data augmentation on cervical abnormality screening, which is comprised of two stages. The first Global Image Generation stage can generate high-resolution normal images full of NILM cells and other natural impurities, which provide visually plausible backgrounds with accurately simulated cell spatial relationships. In Local Cell Editing stage, a certain number of NILM cells selected from the generated normal images can be converted to abnormal cells of different cervical malignant degrees, thus obtaining synthetic abnormal cytopathological images with box annotations. The proposed Normal Image Generator and Abnormal Cell Synthesizer in both stages are built upon pre-trained Stable Diffusion via PEFT techniques for cytopathological image content customization and spatial layout controllability extension, respectively. The utilization of PEFT techniques significantly reduces the computation budget and the need for large-scale data in training reliable generative models. Through explicit manipulation of the numbers, the locations, and the cell types of synthetic abnormal cells, large-scale and diverse annotated abnormal images can be created for augmenting cervical abnormality screening in various clinical scenes. Qualitative and quantitative experiments validate the semantic realism, the flexible controllability, and the data augmentation effectiveness of our proposed synthesis framework.

### 25 Feb 2024 v2

@@ Abstract
Automatic thin-prep cytologic test (TCT) screening can assist pathologists in finding cervical abnormality towards accurate and efficient cervical cancer diagnosis. Current automatic TCT screening systems mostly involve abnormal cervical cell detection, which generally requires large-scale and diverse training data with high-quality annotations to achieve promising performance. Pathological image synthesis is naturally raised to minimize the efforts in data collection and annotation. However, it is challenging to generate realistic large-size cytopathological images while simultaneously synthesizing visually plausible appearances for small-size abnormal cervical cells. In this paper, we propose a two-stage image synthesis framework to create synthetic data for augmenting cervical abnormality screening. In the first Global Image Generation stage, a Normal Image Generator is designed to generate cytopathological images full of normal cervical cells. In the second Local Cell Editing stage, normal cells are randomly selected from the generated images and then are converted to different types of abnormal cells using the proposed Abnormal Cell Synthesizer. Both Normal Image Generator and Abnormal Cell Synthesizer are built upon Stable Diffusion, a pre-trained foundation model for image synthesis, via parameter-efficient fine-tuning methods for customizing cytopathological image contents and extending spatial layout controllability, respectively. Our experiments demonstrate the synthetic image quality, diversity, and controllability of the proposed synthesis framework, and validate its data augmentation effectiveness in enhancing the performance of abnormal cervical cell detection.
@@ I Introduction
Recently, diffusion models have gained enormous popularity due to their remarkable capabilities to generate high-quality images [13, 14]. Learning from large-scale data, text-to-image (T2I) diffusion models, especially Stable Diffusion [15], have demonstrated unprecedented ability in creating a wide variety of image contents that closely resemble the input text descriptions. These T2I diffusion models serve as foundation models for image synthesis, which can be adapted to a wide range of domains and tasks through transfer learning. To unleash their potentials in a computationally efficient way, many studies have explored parameter-efficient fine-tuning (PEFT) methods in pursuit of model customization [16, 17, 18, 19] or additional controllability [20, 21, 22]. The impressive progress has inspired us to employ PEFT techniques to tailor T2I diffusion models for both few-shot synthesis and fine-grained conditioning in cervical cytopathological image synthesis.
In this paper, we propose a novel cytopathological image synthesis framework to augment cervical abnormality screening by adapting T2I foundation models to the pathology domain. The proposed framework can synthesize diverse cytopathological images that incorporate cervical cells of various abnormality degrees as shown in Fig. 1. The synthesis process naturally brings in a bounding box when placing an abnormal cell in the generated image. The synthetic data can thus be used to train a detector for suspicious abnormal cervical cells, which is essential to clinical usage. We empirically set the synthetic image size to be $1024\times 1024$ with a microscope magnification of ×$\times 20$, which strikes a balance between the number of cells in each image and the content of the image that a detector can effectively handle. As illustrated in Fig. 2, the overall framework is decomposed into two stages below:
@@
  • Local Cell Editing. Based on a hypothesis that the neighboring effects between cervical cells are eliminated after oscillation and centrifugation in WSI preparation, we propose to synthesize abnormal cells by individually editing the cell types. Specifically, we employ a simple in-house cell detector to localize the synthetic NILM cells. After randomly selecting a certain number of them, we apply the Abnormal Cell Synthesizer. It can effectively extend the conditioning functionality of the pre-trained Stable Diffusion and translate the selected NILM cells into abnormal cells of user-defined types.
@@
  • We incorporate PEFT techniques in building Normal Image Generator and Abnormal Cell Synthesizer upon Stable Diffusion, the state-of-the-art pre-trained foundation model for image synthesis, thus significantly alleviating the need for large-scale data in training reliable medical generative models.
@@ III Method
In this paper, we propose to disentangle the overall cervical cytopathological image synthesis into two stages, namely Global Image Generation and Local Cell Editing. As demonstrated in Fig. 2, the first stage produces high-resolution normal cytopathological images, and the second stage synthesizes high-fidelity abnormal cell patches. In the next, Section III-A reviews preliminary information about Stable Diffusion, which serves as the pre-trained foundation model in our work. Section III-B and Section III-C present Global Image Generation and Local Cell Editing in detail, respectively.
@@ V Conclusion and Discussion
Cytopathological image synthesis can minimize the efforts in data collection and annotation for cervical abnormality screening. However, it is difficult to generate a large-size cytopathological image while synthesizing visually realistic cellular features for small-size abnormal cells. To overcome this challenge, we propose a novel two-stage cervical cytopathological image synthesis framework based on Stable Diffusion, a foundation model for image synthesis, for data augmentation on cervical abnormality screening. The first Global Image Generation stage can generate high-resolution normal images full of NILM cells and other natural impurities, which provide visually plausible backgrounds with accurately simulated cell spatial relationships. In Local Cell Editing stage, a certain number of NILM cells selected from the generated normal images can be converted to abnormal cells of different cervical malignant degrees, thus obtaining synthetic abnormal cytopathological images with box annotations. The proposed Normal Image Generator and Abnormal Cell Synthesizer in both stages are built upon pre-trained Stable Diffusion via PEFT techniques for cytopathological image content customization and spatial layout controllability extension, respectively. The utilization of PEFT techniques significantly reduces the computation budget and the need for large-scale data in training reliable generative models. Through explicit manipulation of the numbers, the locations, and the cell types of synthetic abnormal cells, large-scale and diverse annotated abnormal images can be created for augmenting cervical abnormality screening in various clinical scenes. Qualitative and quantitative experiments validate the semantic realism, the flexible controllability, and the data augmentation effectiveness of our proposed synthesis framework.